38. Solve the initial value problem x+3y=+*2 + 2xy +T Y() = -2 2xy x2 + 2x2y + 1 Ti y(1) = -2.
Answer #1

38 y+ put y = y = vx. Then -2 = van de V + d +9V - 27th -+2**v + * * *+2x3v11 - 3v = - 20² + 3x²~ +603 v² + 3v. x²+2x², +10 0 S (x +exv++) av = *v + v + = y + y2 + y Ev=4] (5v + 6xv+ a) dhe ( is superand) = 5Vx+ 3v - 3V = 5y + 3y ² - 3 2 2 y u=